The temperature at the surface of the sun is 6000K. Convert this to Celsius, Fahrenheit, and Rankine.

1
Expert's answer
2022-04-27T13:34:31-0400
"6000K=(6000-273)\: C=5727\: C"

"6000K=((6000-273)*1.8+32)\: F=10341\: F"

"6000K=(6000*\frac{9}{5})\: R=10800\: R"